The distance between two successive crests in a wave is 1.25 m, and the source generates 15 crests in 5.0 s. What is the speed of the wave?


1
Expert's answer
2022-01-28T08:14:10-0500

Given:

"\lambda=1.25\:\rm m"

"N=15"

"t=5.0\:\rm s"


The frequency of oscillation

"f=\frac{N}{t}=\frac{15}{5.0}=3.0\:\rm Hz"

Hence, the speed of the wave

"v=f\lambda=3.0*1.25=3.75\:\rm m/s"